3x+y=21
10x+5y=65
: 3x+y=21
10x+5y=65

Answer by stanbon(18743)

3x+y=21
10x+5y=65
---------------
3x+y=21
10x+5y=65
-----------------------
Multiply thru 1st equation by 5:
15x+5y= 105
10x+5y=65
-----------
Subtact 2nd from 1st to solve for "x":
5x = 40
x = 8
--------------
Substitute into 3x+y = 21 to solve for "y":
3*8 + y = 21
y = -3
--------------
Ans: (8,-3)
===================
Cheers,
Stan H.